### v1

# XIFBench: Evaluating Large Language Models on Multilingual Instruction Following

Abstract
Large Language Models (LLMs) have demonstrated remarkable instruction-following capabilities across various applications. However, their performance in multilingual settings remains poorly understood, as existing evaluations lack fine-grained constraint analysis. We introduce XIFBench, a comprehensive constraint-based benchmark for assessing multilingual instruction-following abilities of LLMs, featuring a novel taxonomy of five constraint categories and 465 parallel instructions across six languages spanning different resource levels. To ensure consistent cross-lingual evaluation, we develop a requirement-based protocol that leverages English requirements as semantic anchors. These requirements are then used to validate the translations across languages. Extensive experiments with various LLMs reveal notable variations in instruction-following performance across resource levels, identifying key influencing factors such as constraint categories, instruction complexity, and cultural specificity.  [^1] Our code and data will be released upon acceptance.
XIFBench: Evaluating Large Language Models on
Multilingual Instruction Following
Zhenyu Li1, Kehai Chen1,thanks:   Corresponding author., Yunfei Long2, Xuefeng Bai1, Yaoyin Zhang3, Xuchen Wei3, Juntao Li4, Min Zhang1 1Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China 2School of Computer Science and Electronic Engineering, University of Essex, Colchester, United Kingdom 3School of Computer Science and Technology, Harbin Institute of Technology, Shenzhen, China 4School of Computer Science and Technology, Soochow University, China 23s051032@stu.hit.edu.cn, chenkehai@hit.edu.cn, yl20051@essex.ac.uk
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in various natural language processing (NLP) tasks and real-world applications (Zhao et al., 2023). A key factor behind their widespread adoption is their ability to follow instructions (Ouyang et al., 2022; Wei et al., 2022), allowing them to comprehend and execute user commands while adhering to specified requirements (Zhou et al., 2023b; He et al., 2024b). Consequently, ensuring consistent multilingual instruction-following capability has attracted significant attention in advancing LLMs (Qin et al., 2024a; Shaham et al., 2024).
Figure 1: Discrepancies in LLMs’ instruction following across languages. Given English instruction and Chinese translation, Llama-3.1-8B responses show varying capabilities in the following constraints, as seen in constraint-based evaluation. A concise back-translation (b.t.) of responses is provided for reference.
Despite the promise of multilingual instruction following, significant challenges remain due to disparities in language resources (Figure 1). Recent studies (Shaham et al., 2024; Weber et al., 2024) reveal substantial discrepancies in LLMs’ instruction-following capabilities across high- and low-resource languages. Most existing evaluations rely on simple, direct instructions and holistic assessment methods, such as pairwise comparison (Li et al., 2023) or direct scoring (Zheng et al., 2023), which fail to capture the complexity of real-world instructions and provide limited insights into performance gaps. Constraint-based evaluations have emerged as a promising alternative (Jiang et al., 2024; Qin et al., 2024b), as constraints encapsulate essential requirements that LLMs should satisfy. Constraints not only serve as fundamental components for instruction construction and evaluation but also enable a more fine-grained analysis of instruction-following capabilities in high-resource languages like English and Chinese. However, this leaves a critical gap in understanding how LLMs handle constraints across diverse languages, particularly for medium- and low-resource ones.
Motivated by this challenge, this paper aims to quantify and analyze multilingual instruction-following variations and their causes systematically. To achieve this, we introduce XIFBench, a comprehensive benchmark featuring three key innovations: (1) It is built upon a novel constraint taxonomy covering five categories—Content, Style, Situation, Format, and Numerical—with 21 dimensions carefully selected to ensure intra-category consistency and cross-lingual applicability by excluding language-dependent constraints. This taxonomy enables the automatic construction of instruction sets with balanced constraint categories and difficulty distributions. (2) To ensure consistent evaluation across languages, we develop a requirement-based protocol that decomposes complex instructions into evaluation requirements—English YES/NO questions serving as semantic anchors for cross-lingual assessment. This approach enables fine-grained, interpretable evaluation of instruction-following capabilities. (3) XIFBench provides high-quality parallel instruction sets in six languages, spanning high-resource (English, Chinese), medium-resource (Russian, Arabic), and low-resource (Hindi, Swahili) languages, with validation of requirement equivalence across translations. Extensive experiments with various LLMs reveal notable variations in instruction-following performance across resource levels, identifying key influencing factors such as constraint categories, instruction complexity, and cultural specificity.
2 Related Work
Instruction Following Evaluation
As the fundamental capability of LLMs to align with human intents, InstructGPT (Ouyang et al., 2022) evaluated instruction following through human evaluation, employing direct scoring and binary adherence checks. Later works like AlpacaEval (Li et al., 2023) and MT-Bench (Zheng et al., 2023) utilized LLM judges for pairwise comparison and direct scoring, respectively. To assess LLMs on more complex tasks, WizardLM (Xu et al., 2024) and LIMA (Zhou et al., 2023a) introduced diverse and challenging instructions, although they still relied on holistic evaluation methods.
The advancement of constraint-based evaluation has enabled fine-grained evaluation of instruction following. CELLO (He et al., 2024a) and CoDI-Eval (Chen et al., 2024) advanced this direction by collecting complex instructions and applying rule-based constraint checks. While IFEval (Zhou et al., 2023b) focused solely on rule-verifiable format constraints, FollowBench (Jiang et al., 2024) expanded to multiple constraint types (e.g., content, situation, style) and analyzed the impact of constraint quantity with LLM-based evaluation. InfoBench (Qin et al., 2024b) introduced requirement checklists for precise assessment and has been widely adopted in later studies. CFBench (Zhang et al., 2024a) developed a comprehensive constraint taxonomy from real-world scenarios, and ComplexBench (Wen et al., 2024) further investigated the impact of constraint composition types. However, most studies primarily focus on English and Chinese, leaving a gap in understanding instruction across languages.
Multilingual Instruction Following Evaluation
Evaluating instruction following in multilingual contexts often relies on translated English-centric benchmarks like AlpacaEval, as demonstrated in various instruction tuning and evaluation studies (Zhang et al., 2024b; Li et al., 2024a; Shaham et al., 2024; Weber et al., 2024). Some works introduce language-specific benchmarks, such as Arabic MT-Bench (Boughorbel and Hawasly, 2023) and CIFBench (Li et al., 2024b) for Chinese. Although Aya Evaluation Suite (Singh et al., 2024) provides native annotations and translations across 101 languages, its evaluation approaches remain holistic.
Constraint-based evaluation remains underexplored in multilingual settings. A notable effort in this direction is Multi-IF (He et al., 2024b), which extends IFEval to multi-turn scenarios across multiple languages. However, Multi-IF primarily focuses on format constraints, leaving broader aspects of constraint adherence underexplored. In this paper, we propose XIFBench, an automatically constructed multilingual benchmark with a comprehensive constraint taxonomy for evaluating instruction following across resource levels.
Figure 2: The automated pipeline for constructing XIFBench, consisting of three stages with six steps: Constraint Augmentation (§3.1.2), Requirement Structuring (§3.1.3), and Multilingual Expansion (§3.1.4).
3 XIFBench Dataset
3.1 Dataset Construction
We construct XIFBench by preparing diverse seed instructions, then proceed with three automated stages: (1) augmenting them with constraints, (2) extracting fine-grained evaluation requirements, and (3) expanding the dataset to multiple languages. Figure 2 illustrates the construction workflow.
3.1.1 Instruction Preparation
Seed Instruction Selection
To ensure the diversity and representativeness of XIFBench, we source evaluation instructions from AlpacaEval (Li et al., 2023), WizardLM (Xu et al., 2024), and LIMA (Zhou et al., 2023a), which are widely used benchmarks for assessing multilingual instruction following capabilities (Zhang et al., 2024b; Li et al., 2024a; Shaham et al., 2024; Weber et al., 2024). To prevent data leakage, we include only evaluation sets and omit all training data. To promote task diversity, we perform hierarchical clustering on these instructions, resulting in 131 distinct clusters. We then select one representative instruction from each cluster, ensuring a diverse and comprehensive instruction set. See Appendix A.1.1 for details.
Instruction Filtering and Annotation
While existing benchmarks provide a useful starting point, not all instructions are suitable for multilingual evaluation or further augmentation. Thus, we manually filter out ambiguous, overly difficult, or language-dependent instructions (e.g., capitalized response, alliteration), yielding a refined set of 106 instructions, named the Easy Set. Each instruction is further annotated with cultural accessibility labels to indicate culturally specific references (e.g., Jesus Christ, YouTube) that may not be universally understood across languages. Additionally, we decompose each instruction into core instruction and input materials to facilitate subsequent automated construction. See Appendix A.1.2 for details.
3.1.2 Constraint Augmentation
Constraint Taxonomy
To systematically evaluate the impact of different constraint categories on multilingual instruction following, we propose a novel taxonomy comprising 5 categories and 21 dimensions, inspired by prior constraint-based evaluations (Jiang et al., 2024; Qin et al., 2024b; Zhang et al., 2024a). As illustrated in Figure 2, each constraint follows the structure category - dimension - specification, such as content - include content - discuss scientific research, guiding LLMs to include scientific research in their responses.
Our taxonomy covers five major categories: (1) Content, which specifies what information should be included or excluded; (2) Style, which defines the tone and writing manners; (3) Situation, which describes contextual settings such as assumed roles or environments; (4) Format, which outlines structural requirements for responses; and (5) Numerical, which involves quantitative factors like length or item counts. To ensure cross-lingual applicability, we intentionally exclude Linguistic constraints that depend on language-specific properties (e.g., uppercase, alliteration, word count), which were a focus in prior English- and Chinese-centric taxonomies. See Appendix A.2.1 for details.
Constraint Brainstorming
For each core instruction, we prompt GPT-4o (OpenAI, 2024) [^2] We use gpt-4o-2024-08-06 for all constructions. to generate multiple independent constraints across all 5 categories and 21 dimensions, ensuring that each dimension is adequately covered. This process results in a balanced constraint set while maximizing the variety of applicable constraints for each instruction. Additionally, it allows precise control over the types of constraints introduced in subsequent steps. See Appendix A.2.2 for details.
Constraint Combination
To balance instruction difficulty and diversity, we sample constraints from the brainstormed set to construct instructions with 1 to 5 additional constraints. The sampling algorithm prioritizes underrepresented categories for even distribution. We then prompt GPT-4o to integrate the selected constraints into fluent and natural instruction. This process yields the Hard Set, containing 530 combined instructions. See Appendix A.2.3 for details.
Instruction Validation
To ensure instruction quality, two human annotators independently assess each instruction for clarity, cultural accessibility, and linguistic dependency (i.e., whether it involves linguistic constraints). Instructions with defects are removed along with all their variations in both the Easy Set and Hard Set to maintain consistency. This validation yields 93 high-quality instructions in the Easy Set and 465 in the Hard Set. See Appendix A.2.4 for details.
3.1.3 Requirement Structuring
Requirement Decomposition
To enable fine-grained evaluation, we decompose instructions into atomic requirements, following prior works Qin et al. (2024b); Zhang et al. (2024a); Wen et al. (2024). Moreover, we automatically complete this process by prompting GPT-4o to extract evaluation requirements from the instructions in both Easy Set and Hard Set. As shown in Figure 2, each requirement is a binary (YES/NO) question (e.g., Is each section limited to 2 sentences?), ensuring a precise and interpretable assessment. See Appendix A.3.1 for details.
Requirement Categorization
Since the initial core instructions contain inherent constraints, and constraints from the combination step may be rephrased differently, we need to map the extracted evaluation requirements back to their original categories. To achieve this, we prompt GPT-4o to classify each requirement into the most suitable category and dimension within our predefined taxonomy. This automated process ensures consistency and facilitates a detailed analysis of the adherence to constraints in languages and models. See Appendix A.3.2 for details.
Requirement Assessment
To verify the accuracy of requirement extraction and categorization, we sample 10% of instructions (10 from the Easy Set, 50 from the Hard Set) and ask two annotators to independently evaluate the extracted requirements on four dimensions: (1) Explicitness: The requirement must be explicitly stated in the instruction. (2) Completeness: It must cover all explicitly stated constraints. (3) Atomicity: It should be atomic, addressing a single aspect. (4) Categorization: It should be assigned to the correct category. Our evaluation confirms its high reliability. See Appendix A.3.3 for details.
3.1.4 Multilingual Expansion
Language Selection
To ensure broad multilingual coverage while maintaining feasibility, we select six representative languages—English, Chinese, Russian, Arabic, Hindi, and Swahili—based on prior multilingual instruction tuning works (Shaham et al., 2024; Li et al., 2024a). These languages span different resource levels: high (English, Chinese), medium (Russian, Arabic), and low (Hindi, Swahili), following Joshi et al. (2020). This selection enables a systematic analysis of language resource impact on instruction following.
Parallel Translation
We translate the core instruction of each English instruction in Easy Set and the combined instruction in Hard Set into five other languages using Google Translate, while keeping the input materials unchanged. This ensures that both code snippets and structured formats remain consistent across languages.
Translation Validation
While automatic translation is generally reliable, subtle semantic shifts may still affect constraint-based evaluation. As shown in Figure 2, we use GPT-4o and Google Translate’s back-translation to verify requirement equivalence across languages, ensuring each translated instruction segment aligns with the original English version. Our analysis reveals that only 0.29% of evaluation requirements exhibit inconsistencies across languages, demonstrating sufficient translation quality for multilingual evaluation. See Appendix A.4 for details.
3.2 Dataset Statistics
Dataset #Inst. #Req. Avg. Tokens Avg. Req.
Easy Set 93 269 36.3 $\pm$ 53.0 2.9 $\pm$ 1.1
Hard Set 465 1395 69.4 $\pm$ 56.0 5.01 $\pm$ 1.7
Overall 558 1664 48.1 $\pm$ 22.4 4.66 $\pm$ 1.8
Table 1: Dataset statistics of XIFBench, including instance count, evaluation requirements count, and average token lengths.
Table 1 summarizes XIFBench, which contains 558 instructions and 1,664 evaluation requirements. Each instruction averages 48.1 ± 22.4 tokens, including core instructions and optional input materials. To enable multilingual evaluation, instructions are translated into five languages (Chinese, Russian, Arabic, Hindi, and Swahili), yielding 3,348 instances. The evaluation requirements remain unchanged across languages for consistent evaluation. See Appendix B for details.
3.3 Evaluation Protocol
Evaluation Framework
We adopt the LLM-based evaluation framework from InfoBench (Qin et al., 2024b), using predefined evaluation requirements to assess how well the responses follow instructions. For each translated instruction and response, we prompt an LLM to generate concise observations and a binary YES/NO decision for each requirement. To ensure cross-linguistic comparability, we provide the original English instruction, evaluation requirements, and metadata (categories and dimensions), keeping them identical across languages as semantic anchors for consistent evaluation. English responses are evaluated on the original English instructions and requirements. See Appendix C for details.
Evaluation Metrics
We define two primary metrics: Requirement Following Rate (RFR) and Instruction Following Rate (IFR). RFR measures the percentage of evaluation requirements correctly satisfied across all instructions, offering a fine-grained view of adherence to specific constraints. IFR quantifies the percentage of instructions where all requirements are met, providing a stricter assessment of overall compliance. These metrics capture LLMs’ ability to follow instructions across constraints and complexities.
Metrics Formulation
Let $\mathcal{I}^{(l)}$ be the set of instructions in language $l\in\{\text{en},\ldots,\text{sw}\}$, each instruction $i^{(l)}\in\mathcal{I}^{(l)}$ has an associated set of evaluation requirements $\mathcal{R}_{i}$, identical across languages. Given an LLM-generated response $o^{(l)}_{i}$ to instruction $i^{(l)}$, the adherence to a requirement $r\in\mathcal{R}_{i}$ is denoted as $e^{(l)}_{i,r}$, a binary value: $1$ if satisfied, and 00 otherwise. The LLM-based evaluation function $\mathcal{E}$ determines requirement adherence:
$e^{(l)}_{i,r}=\mathcal{E}(i^{(l)},o^{(l)}_{i},i^{(\mathrm{en})},r),$ (1)
where $\mathcal{E}$ returns $1$ if the requirement is met, 00 otherwise. English instructions are simplified to:
$e^{(\mathrm{en})}_{i,r}=\mathcal{E}(i^{(\mathrm{en})},o^{(\mathrm{en})}_{i},r).$ (2)
The RFR for language $l$ is the proportion of satisfied requirements over all instructions:
$\text{RFR}^{(l)}=\frac{\sum_{i^{(l)}\in\mathcal{I}^{(l)}}\sum_{r\in\mathcal{R}_{i}}e^{(l)}_{i,r}}{\sum_{i^{(l)}\in\mathcal{I}^{(l)}}|\mathcal{R}_{i}|}.$ (3)
The IFR is the proportion of instructions where all requirements are satisfied:
$\text{IFR}^{(l)}=\frac{1}{|\mathcal{I}^{(l)}|}\sum_{i^{(l)}\in\mathcal{I}^{(l)}}\prod_{r\in\mathcal{R}_{i}}e^{(l)}_{i,r}.$ (4)
4 Experiments
4.1 Automatic Evaluation
Metric Requirement Following Rate (RFR) Instruction Following Rate (IFR)
Language En Zh Ru Ar Hi Sw Avg. En Zh Ru Ar Hi Sw Avg.
Closed-Source Language Models
GPT-4o 93.6 92.5 92.7 90.8 92.8 90.8 92.2 76.9 73.3 74.2 69.2 73.8 65.6 72.2
Gemini-2.0-Flash 93.3 93.0 93.0 91.9 92.0 89.5 92.1 78.1 76.7 77.0 75.8 71.7 69.2 74.7
Claude-3.5-Sonnet 89.1 81.3 84.6 75.9 76.1 74.5 80.2 66.1 53.0 61.6 46.1 43.5 40.1 51.8
Open-Source Language Models
Llama-3.1-70B 91.7 83.4 87.3 76.4 80.9 73.4 82.2 70.9 48.9 58.2 40.2 42.7 34.8 49.3
Qwen-2.5-72B 90.5 89.1 89.2 85.4 82.3 40.9 79.6 67.7 63.3 64.9 57.2 51.3 10.4 52.5
Qwen-2.5-14B 89.7 88.5 87.8 82.1 70.0 23.5 73.6 64.7 61.1 60.6 51.4 33.9 5.0 46.1
Glm-4-9B 86.4 87.2 85.0 78.6 71.3 25.6 72.4 56.5 60.3 54.2 43.2 34.6 4.1 42.2
Llama-3.1-8B 87.6 79.1 79.5 63.2 66.8 38.6 69.1 58.9 42.8 44.9 25.3 27.2 9.7 34.8
Qwen-2.5-7B 87.8 87.4 83.3 77.1 59.6 10.0 67.6 59.9 57.3 52.3 41.0 21.1 1.1 38.8
Table 2: Following rates (%) of different models on XIFBench. The highest rates among closed-source models are in bold, while those among open-source models are underlined. Avg. represents the average rate across all languages.
We employ GPT-4o [^3] We use gpt-4o-2024-08-06 for all evaluations. as the evaluation judge to assess 9 LLMs across two categories: (1) Closed-source LLMs: GPT-4o [^4] We assess gpt-4o-2024-11-20 to avoid self-evaluation., Gemini-2.0-Flash (Google, 2025), and Claude-3.5-Sonnet (Anthropic, 2024); (2) Open-source LLMs: Llama-3.1-70B, Llama-3.1-8B (Meta, 2024), Qwen-2.5-72B, Qwen-2.5-14B, Qwen-2.5-7B (Qwen, 2025), and GLM-4-9B-Chat (GLM, 2024). We access closed-source models via their official APIs, while open-source models run locally using vLLM (Kwon et al., 2023) for efficient inference. We use greedy decoding with a maximum token limit of 4096 to ensure determinism and prevent truncation, keeping all other hyperparameters at default values. LLMs are tested on both sets. We report RFR and IFR scores for each LLM.
Table 2 presents the main results of the automatic evaluation. From a language resource perspective, we observe that: (1) Performance correlates with resource levels. High-resource languages (e.g., English, Chinese) exhibit stable performance, medium-resource ones (e.g., Russian, Arabic) show greater variance, and low-resource languages (e.g., Hindi, Swahili) experience steep IFR declines, sometimes nearing zero. (2) The RFR-IFR gap is most pronounced in low-resource languages. Models achieve reasonable RFR scores but struggle with IFR, indicating difficulties in full instruction adherence. For example, in Hindi, several models exceed 80% in RFR but fall below 40% in IFR. (3) Instruction adherence remains a challenge across all languages. The persistent RFR-IFR gap suggests that limitations extend beyond data scarcity—models can follow individual constraints but struggle with holistic instruction execution.
From a model capability perspective, we observe in Table 2 that: (1) Closed-source models excel in RFR but still face IFR challenges. Despite the strong performance, models like GPT-4o and Gemini-2.0-Flash achieve RFR above 90% but IFR only around 70%, revealing limitations in full adherence. (2) Open-source models exhibit sharper IFR declines. While their RFR remains stable in high-resource languages, it degrades significantly in lower-resource settings, indicating weaker cross-lingual generalization. (3) Larger models generally outperform smaller ones, though exceptions (e.g., Qwen-2.5-14B in Swahili) likely stem from data composition rather than scaling limitations.
4.2 Agreement Evaluation
Methods En Zh Ru Ar Hi Sw Avg.
DS 70.0 56.7 56.7 55.0 56.7 70.0 60.9
Ours 95.8 96.8 95.1 93.0 95.4 92.3 94.7
Table 3: Agreement rates between GPT-4o and human annotations across different languages (%) within DS and our evaluation protocol. bold means best results.
Figure 3: Cross-lingual RFR performance comparison across constraint categories for three representative models. Each radar chart illustrates the RFR scores across different languages within each constraint category.
Since LLM-based evaluators vary in agreement with humans across languages Hada et al. (2024), we assess the proposed protocol’s reliability through agreement studies. We randomly sample 360 (instruction, response) pairs from six languages, covering both the Easy Set and Hard Set. These pairs are generated by three LLMs with varying scales and sources: Gemini-2.0-Flash, Qwen-2.5-72B, and Glm-4-9B. Other settings follow Section 4.1. We compare our protocol with Direct Scoring (DS) Zheng et al. (2023) whose responses are rated from 1 to 10. Following Jiang et al. (2024); Qin et al. (2024b), responses scoring above 5 are considered to meet all evaluation requirements. Two human annotators assess whether GPT-4o’s evaluation aligns with human judgment for each ground-truth evaluation requirement. GPT-4o also performs Direct scoring for comparison. Finally, we compute agreement rates between GPT-4o and human annotations to assess our protocol’s consistency. See Appendix D.2 for details.
As shown in Table 3, our evaluation protocol outperforms DS across languages, with a slight drop in low-resource languages. The unexpectedly high agreement rate of DS in Swahili does not indicate greater reliability but results from two of the three selected models performing poorly—often scoring below 5 and classified as "not all requirements met"—aligning with human annotations and incidentally boosting agreement.
5 Analysis
5.1 How Do Constraint Categories Affect Cross-lingual Performance?
Figure 4: Cross-lingual IFR performance comparison across instruction complexity levels for three representative models. Each group shows IFR scores across languages per constraint count.
Figure 5: Cross-lingual RFR and IFR performance comparison of culturally universal (CUI) and specific (CSI) instructions across three models. Each group presents RFR and IFR scores per instruction type across languages.
We only use the Hard Set for its balanced constraint distribution and focus on RFR to align with the granularity of constraint analysis. We analyze three models spanning different capabilities: Gemini-2.0-Flash, Qwen-2.5-72B, and Glm-4-9B. Other settings follow Section 4.1.
As illustrated in Figure 3, we find that higher-capacity models exhibit more robust cross-lingual constraint adherence. (1) For high-capacity models like Gemini-2.0-Flash, the radar plots form nearly regular polygons across languages, indicating stable adherence with minimal cross-lingual variation. (2) In contrast, mid- and low-capacity models (Qwen-2.5-72B, Glm-4-9B) exhibit progressive degradation in polygon size and regularity as language resources decrease. This degradation is most pronounced for style and situation constraints, suggesting these constraints are more sensitive to language-specific factors.
Further examining constraint categories, we observe that some constraints exhibit universal properties, while others are language-dependent. (1) Format and numerical constraints are more resilient to language variations, as indicated by denser polygon edges (i.e., more consistent performance). This suggests they rely on universal linguistic properties. (2) In contrast, style, situation, and content constraints are more sensitive to language resources, with sparser polygon edges as resources decline. Style and situation constraints degrade the most, reinforcing their strong dependence on language-specific properties, while content constraints show moderate degradation.
5.2 How Does Instruction Complexity Impact Cross-lingual Performance?
We use both the Easy Set and Hard Set to obtain instructions ranging from no additional constraints to five constraints. IFR is used to align with instruction-level analysis. Model selection and other experimental settings follow Section 5.1.
As shown in Figure 4, we find that stronger LLMs maintain more consistent performance across instruction complexity levels. (1) Gemini-2.0-Flash demonstrates stable performance across all language resource levels without additional constraints and exhibits relatively smooth degradation as constraints increase, maintaining acceptable IFR even with five constraints. (2) The mid- and low-capacity models (Qwen-2.5-72B, Glm-4-9B) not only perform slightly worse but also degrade more steeply across languages, dropping to unacceptable IFR levels with five constraints in Swahili.
Further analysis reveals that higher-resource languages experience lower performance degradation under complex instructions. (1) High-resource languages (English, Chinese) show steady degradation while maintaining acceptable absolute performance. (2) In contrast, mid- and low-resource languages exhibit slightly sharper degradation, resulting in substantial IFR gaps across resource levels as complexity increases.
5.3 To What Extent Does Cultural Specificity Influence Instruction Following?
We use the Easy Set and Hard Set to compare LLM performance on culturally universal and specific instructions (Section 3.1). Both RFR and IFR are included for a comprehensive view of instruction following. Model selection and other settings follow Section 5.1.
As shown in Figure 5, we observe several key findings: (1) As language resources decrease, both RFR and IFR decline for culturally universal and specific instructions. The degradation rate is slightly steeper for culturally specific instructions but not substantially. (2) More capable models like Gemini-2.0-Flash show smaller performance gaps between culturally universal and specific instructions, while smaller models like GLM-4-9B degrade more on culturally specific instructions. (3) IFR is more sensitive to cultural specificity than RFR, especially as language resources or model capacity decrease, since failures in cultural constraints hinder complete instruction following. These findings suggest that despite pre-training exposure to cultural content, cultural specificity still impacts instruction following, particularly for smaller models and low-resource languages.
6 Conclusion
In this paper, we introduced XIFBench, a benchmark for evaluating the multilingual instruction-following capabilities of LLMs. It features a novel constraint taxonomy, a requirement-based evaluation protocol, and high-quality parallel instruction sets across six languages. Through extensive experiments, we revealed notable variations in LLM performance across resource levels, constraint categories, and instruction complexities. Our findings reveal that LLMs struggle with complex constraints and low-resource languages, highlighting the need for better multilingual training and constraint-aware optimization.
Limitations
While XIFBench provides a comprehensive evaluation framework, certain aspects remain constrained by our design choices: (1) Our benchmark focuses on general-purpose LLMs, excluding reasoning-oriented models (e.g., o1, DeepSeek-R1) due to budget and API constraints. These models, which often infer missing constraints or explore alternative solutions, may exhibit different patterns in constraint adherence, a factor not covered in our current study. (2) As our benchmark relies on parallel translations, cultural accessibility annotations are primarily based on English contexts, which may not fully capture the diversity of cultural factors influencing instruction adherence. Incorporating native speaker annotations could provide a more nuanced perspective. (3) While we identify key factors affecting multilingual instruction following, leveraging these insights to enhance LLMs remains an open challenge. Future research could explore how constraint-based evaluation can inform training or adaptation to improve instruction adherence.
Appendix A Data Construction Details
In this part, we provide detailed information on the construction of XIFBench, including instruction preparation, constraint augmentation, requirement structuring, and multilingual expansion.
A.1 Instruction Preparation Details
A.1.1 Seed Instruction Selection
As described in Section 3.1.1, we source instructions from AlpacaEval, WizardLM, and LIMA via their official repositories on Hugging Face and GitHub. Since WizardLM and LIMA primarily contribute instruction-tuning datasets along with evaluation sets, we use only the evaluation sets to construct our benchmark, thereby avoiding potential data leakage.
For hierarchical clustering, we utilize the best-performing model, all-mpnet-base-v2, in Sentence Transformers [^5] https://www.sbert.net/, to encode the instructions and compute the Euclidean distances between them. We then apply hierarchical clustering to group similar instructions. To determine an appropriate number of clusters, we iterate over various threshold distances and select a value that ensures both diversity and sufficient seed instructions. Ultimately, we set the threshold distance to 1.8, resulting in 131 clusters. To maximize diversity, we randomly select a single instruction from each cluster.
A.1.2 Instruction Filtering and Annotation
Instruction Validation Task
Task Introduction
Your task is to validate each instruction against specific rules given in the criteria section. We will use instructions that meet these criteria as seed instructions to create a multilingual instruction-following benchmark for Large Language Models (LLMs). We will complexify these seed instructions by adding constraints to create a hard set of instructions for LLMs. We will translate these instructions into different languages such as English, Chinese, Russian, Arabic, Hindi, Swahili, and more.
Input & Output Explanation
{{INPUT_OUTPUT_EXPLAINATION}}
Validation Criteria
Please check each instruction against the following criteria step-by-step:
Rule 1: Is the instruction clear to understand? [YES/NO]
Definition: The instruction should be clear to understand, avoiding ambiguity or complexity. This ensures the instruction can be understood by LLMs with no confusion. Positive Example: "What is the meaning of life?"
Accepted Reason: The instruction is clear to understand. Negative Example: "Summarize this website LOTTADIGITAL.COM"
Rejected Reason: The instruction is ambiguous due to the lack of context of the website. {{…}}
Rule 2: Does the instruction maintain linguistic independence across different languages? [YES/NO]
Definition: The instruction should be language-neutral, avoiding any features specific to particular languages. This ensures the instruction can be applied across different languages without linguistic barriers. Linguistic features are language-specific elements such as unique alphabets, wordplay, puns, grammatical structures, or punctuation requirements that may not be universally applicable. Positive Example: "Write a paragraph about black holes."
Accepted Reason: The instruction is linguistically independent and can be applied across various languages. Negative Example: "Write a paragraph about black holes using all vowels (A, E, I, O, U)."
Rejected Reason: The instruction relies on English-specific features (vowels), making it unsuitable for other language systems. {{…}}
Rule 3: Does the instruction maintain cultural accessibility across different languages? [YES/NO]
Definition: The instruction should focus on universal concepts while avoiding culturally specific references that may not be globally accessible or understood. This ensures the instruction can be applied across different languages without cultural barriers. Cultural references are elements that either have limited geographic/linguistic spread (such as regional platforms, products, customs, or policies) or require specific cultural background knowledge (such as historical, religious, or traditional contexts). These elements may not be universally accessible or meaningful across different cultural settings. Positive Example: "Why did humans evolve to believe in gods?"
Accepted Reason: The instruction discusses a universal human phenomenon. Negative Example: "Why did humans evolve to believe in Jesus Christ?"
Rejected Reason: The instruction references a specific religious figure, making it less universally accessible. {{…}}
Rule 4: Does the instruction contain equal to or fewer than 5 explicit constraints? [YES/NO]
Definition: The instruction should contain $\leq 5$ explicit constraints. This ensures the instruction is not too complex after adding constraints. Constraints are specific requirements that LLMs must follow in their responses. These are categorized into five types: content, style, situation, format, and numerical constraints. Positive Example: "Write a short story about a detective solving a mystery."
Accepted Reason: The instruction contains 2 constraints: format (short story) and content (detective solving mystery). Negative Example: "Write a short story about a detective solving a mystery. Include a twist at the end. Use descriptive language. Make sure the characters are well-developed. Restrict the story to 500 words."
Rejected Reason: The instruction contains more than 5 constraints across content (detective solving mystery, well-developed characters), style (twist at the end, descriptive language), format (short story), and numerical (500 words). {{…}}
Rule 5: Does the instruction contain equal to or fewer than 4 content constraints? [YES/NO]
Definition: The instruction should contain $\leq 4$ content constraints. This ensures the instruction is balanced and not overly focused on content aspects after adding constraints. Content constraints specify the thematic or topical focus of a response, detailing what information should be included or excluded to guide the depth and scope of the content. Positive Example: "Recommend three science fiction books suitable for teenagers."
Accepted Reason: The instruction contains 2 content constraints: science fiction genre and teenage suitability. Negative Example: "Recommend three science fiction books suitable for teenagers. Include a brief summary of each book. Provide the publication year for each book. Mention the author of each book."
Rejected Reason: The instruction contains more than 4 content constraints including the genre, audience, summaries, dates, and authors. {{…}}
Task Workflow
1. Read the prompt field.
2. Evaluate the instruction against each rule.
3. If feeling uncertain, you can (recommended from most to least encouraged):
    - Use web search or external LLMs to understand the instruction better.
    - Test any LLMs with the instruction to see if it produces the expected output.
4. Fill in the human_annotation_tags fields based on your evaluation.
5. Add any additional comments or notes in the notes field if necessary.
Table 4: Human Annotation Guidelines for Instruction Validation Task
Instruction Decomposition Task
Task Introduction
Your task is to decompose each given instruction into input_material and core_instruction.
We will use these decomposed core instructions to create a multilingual instruction-following benchmark for Large Language Models (LLMs). When evaluating LLMs’ performance on this benchmark, we will also provide the input materials alongside their corresponding instructions.
Input Explanation
{{INPUT_EXPLANATION}}
Output Explanation
You need to decompose the prompt into two components:
- input_material: The source content or reference materials that require processing or analysis as part of the task. These materials contain no directives or requirements.
- core_instruction: The primary directive or task to be completed, including task requirements, scenarios, formats, styles, and other relevant information.
Task Guidelines
After reading the full instruction in the prompt field, separate it into core_instruction and input_materials by copying the exact text without modifications. Consider the following guidelines:
1. The core_instruction should retain its complete semantic meaning even when the input materials are removed.
2. input_materials are typically independent components that can be separated from the instruction without altering its overall meaning.
[…]
4. If there are no clear input_materials, consider the entire text as the core instruction.
Decomposition Examples
Example 1
Given Instruction:
"What is the meaning of life?"
Decomposition:
- Input Materials: None
- Core Instruction: "What is the meaning of life?"
Explanation: This instruction contains no separate input materials. Therefore, the entire text is considered the core instruction.
Example 2
Given Instruction:
[…]
Decomposition:
- Input Materials: "Programming for Everybody (Getting Started with Python)"
- Core Instruction: "Design a syllabus for the given course. Students should be given a list of the chapters with brief explanations of each chapter’s purpose."
Explanation: The course title serves as input material, separated from the core instruction by two newline characters.
Table 5: Human Annotation Guidelines for Instruction Decomposition Task
In this section, we provide detailed information on the filtering and annotation process for the selected seed instructions, as described in Section 3.1.1. The authors select data by following the annotation guidelines in Table 4. Annotators are instructed to assign binary true/false values to five dimensions specified in the guideline, with optional comments.
The clarity dimension measures the understandability of the instruction, while linguistic independence ensures the instruction’s language-neutral properties. Cultural accessibility is also annotated to assess dependence on specific references that may not be globally accessible or understandable. Additionally, we impose restrictions of fewer than five explicit constraints and four content constraints to filter out instructions that might become overly complex after augmentation. This process results in an Easy Set containing 106 instructions.
The annotation guidelines for the decomposition of core instructions and input materials are provided in Table 5. Here, core instructions refer to the essential tasks described in the instructions, while input materials serve as supporting information for these tasks. This annotation process is also conducted by the authors.
A.2 Constraint Augmentation Details
A.2.1 Constraint Taxonomy
Category Dimension Specification Examples
Content Include Content Specify themes, concepts, or information that should be present. Discuss the scientific research related to the 5-second rule.
Exclude Content Specify themes, concepts, or information that should be omitted. Avoid mentioning urban legends not related to the 5-second rule.
Narrow Scope Define specific boundaries for the topic. Focus solely on bacterial transfer rates on different surfaces.
Expand Scope Broaden the scope of discussion. Explore cultural perceptions of food safety beyond the 5-second rule.
Connect Elements Establish relationships between concepts or ideas. Link the concept of the 5-second rule to food safety regulations
Compare Elements Analyze similarities and differences between concepts or ideas. Analyze the differences in contamination risk between wet and dry foods.
Analyze Elements Break down and examine parts or aspects of a topic. Examine the role of surface type in bacterial transfer.
Provide Examples Suggest real-world or theoretical instances. Include studies that have tested the 5-second rule in laboratory settings.
Style Writing Style Define the style or genre to write in or express. Use an academic writing style.
Establish Tone Set the tone or attitude of the response. Maintain a neutral tone throughout.
Express Emotion Convey emotions in the response. Convey a sense of skepticism regarding the validity of the rule.
Use Rhetoric Employ rhetorical techniques. Employ metaphors to explain bacterial movement.
Situation Create Situation Establish a specific context or setting. Set the discussion in a high school science class.
Assume Role Assign a particular role or perspective. Assume the perspective of a parent explaining to their child.
Specify Timeline Specify temporal context or chronological parameters. Discuss the 5-second rule in the context of the 1990s.
Establish Purpose Clarify the reason or goal why the instruction is given. The goal is to debunk common food safety myths.
Format Output Format Define the response format or pattern. Present the information in a FAQ format.
Create Hierarchy Establish a hierarchical order for presenting content. Organize information by level of scientific evidence.
Structure Layout Determine how information is organized. Divide the content into sections with clear headings.
Apply Template Use a specific template or structure. Use a case study format to explore the topic.
Numerical Limit Length Restrict the response length in terms of paragraphs or sentences. Restrict to two sentences per section.
Specify Quantities Define the required number of items. Include exactly three scientific studies in the discussion.
Define Ranges Set numerical ranges for elements. Set a timeframe of 5-10 years for study references.
Table 6: XIFBench’s Constraint Taxonomy and Examples
The detailed constraint taxonomy from Section 3.1.2 is presented in Table 6. The examples provided are brainstormed constraints for the instruction Does the 5-second rule exist?, as illustrated in Figure 2.
A.2.2 Constraint Brainstorming
Constraint Brainstorming Task
Your task is to brainstorm potential constraints for the “given instruction”. The resulting “brainstormed constraints” will be sampled and combined with the “given instruction” later.
The goal is to produce a diverse range of constraints that naturally increase the difficulty of the “given instruction”, without relying on language-specific features that don’t translate well.
Constraint Definitions and Categories
{{CONSTRAINT_TAXONOMY}}
Brainstorming Requirements
  • Generate {brainstormed_constraint_count} diverse constraints for each category.
  • Make each constraint focus on a single, specific intention or requirement.
  • Ensure all constraints are independent and combinable with each other.
  • Never provide specific examples or solutions in the constraints.
  • Avoid using language-specific constraints which may not apply universally across languages. For example:
     • When using rhetoric, avoid alliteration, puns, wordplay, and similar devices.
     • When specifying output format, avoid uppercase letters, punctuation marks, and similar elements.
     • When setting length parameters, avoid word counts, character counts, and similar restrictions.
Output Format
Provide the “brainstormed constraints” only using bullet points for each constraint. For example:
### Content Constraints
- <constraint direction>: <constraint specification>
- Include Content: xxx
Task Inputs & Outputs
Here’s the “given instruction”. Let’s brainstorm potential constraints!
### Given Instruction
{given_instruction}
### Brainstormed Constraints
Table 7: Prompt for Constraint Brainstorming Step
The prompt used for constraint brainstorming, as described in Section 3.1.2, is presented in Table 7. We provide both the core instruction and input materials to GPT-4o, instructing it to generate 12 constraints for each category. We set the maximum generation length to 2048 tokens and use a temperature of 0 to ensure stable results.
A.2.3 Constraint Combination
Constraint Combination Task
Your task is to combine the "given instruction" with all "brainstormed constraints". The resulting "combined instruction" serves as a request or query for Large Language Models (LLMs).
The goal is to enhance the complexity of the "given instruction" by integrating the provided "brainstormed constraints".
Combination Requirements
1. Integrate all "brainstormed constraints" reasonably into the "given instruction" without adding or omitting any constraints.
2. Avoid introducing any new details or requirements that are not present in the original "brainstormed constraints".
3. Express each constraint from the "brainstormed constraints" while maintaining its original category’s intent.
4. Preserve the essential information from the "given instruction".
5. Craft a natural and flowing "combined instruction" by:
 - Separate meta-constraint from each main constraint and combine them when appropriate for clarity.
 - Arrange constraints order to enhance readability, such as from general to specific.
 - Using guiding words (like "please", "ensure") and connecting words (like "while", "also") to enhance flow.
Output Format
Provide the "combined instruction" only. For example:
<combined instruction>
Examples
### Given Instruction
Write a code block in Markdown containing an example of a code block in Markdown. Don’t forget those quadruple backticks.
### Brainstormed Constraints
- situation constraint - establish purpose: aim to provide a guide for using markdown in collaborative document creation.
- format constraint - specify output format/pattern: use bold text to highlight markdown syntax within the explanation.
- content constraint - compare elements: discuss differences between inline code syntax and block code syntax in markdown.
### Combined Instruction
As a guide for collaborative Markdown documentation, please create a code block in Markdown that contains an example of a code block in Markdown. Don’t forget those quadruple backticks. Also, please explain the differences between inline code syntax and block code syntax in Markdown, using bold text to highlight all Markdown syntax elements in your explanation.
Task Inputs & Outputs
Here are the "given instruction" and the "brainstormed constraints". Let’s combine them and write a new "combined instruction"!
## Given Instruction
{given_instruction}
## Brainstormed Constraints
{brainstormed_constraints}
## Combined Instruction
Table 8: Prompt for Constraint Combination Step
The prompt used for constraint combination, as described in Section 3.1.2, is shown in Table 8. For each instruction, we sample five times, incrementally selecting 1 to 5 constraints. To ensure category diversity and balance, only unsampled categories are considered at each step. Given a core instruction and the sampled constraints, we prompt GPT-4o to generate five distinct combined instructions using these sampled constraints. The maximum generation length is set to 512 tokens, and a temperature of 0 is used to ensure stable results.
A.2.4 Instruction Valiadation
As described in Section 3.1.2, we conducted a manual annotation to evaluate the quality of the combined instructions in the Hard Set. The annotation process followed the guidelines in Table 4, except for the two constraint-related criteria. Two annotators independently assessed all instructions based on three dimensions: clarity, linguistic independence, and cultural accessibility. To measure inter-annotator agreement, we randomly selected 10% of the Hard Set (50 instances) and included them in both annotators’ assignments. Disagreements in this subset were resolved conservatively: if either annotator marked an instruction as defective or culturally specific, it was classified as such. Both annotators were graduate students in computer science and received prior training on the annotation guidelines. We report the proportion of instructions identified as clear, linguistically independent, and culturally universal, along with inter-annotator agreement scores.
Criteria Ratio (%) Agreement (%)
Clarity 96.6 96.0
Linguistic Independence 100.0 100.0
Cultural Accessibility 70.56 80.0
Table 9: Evaluation results for various criteria, showing the Ratio of instructions meeting each criterion and the Agreement (%) between evaluators.
As shown in Table 9, nearly all instructions are clear and linguistically independent, demonstrating the high quality of the dataset and the reliability of the automatic construction workflow. The strong inter-annotator agreement further supports the consistency of the annotation process. However, the agreement on cultural accessibility is lower, likely due to the subjective nature of this criterion.
Following the annotation, we removed all unclear and linguistically dependent instructions, along with their corresponding variants in both the Easy Set and the Hard Set, ensuring that each instruction retains all variants across 0 to 5 additional constraints. This process also affected a small number of culturally specific instructions. As a result, the proportion of culturally universal instructions in the Hard Set increased slightly from 70.56% to 70.75%, with a total of 329 culturally universal instructions remaining.
A.3 Requirement Structuring Details
A.3.1 Requirement Decomposition
Requirement Decomposition Task
Your task is to decompose the "given instruction" into "evaluation requirements".
The resulting "evaluation requirements" will be used to comprehensively and granularly assess whether a response adheres to the instruction.
Decomposition Rules
- Extract requirements only from the explicit wording of the instruction.
- Each requirement must correspond to a specific part of the instruction.
- Do not infer or add requirements that are not directly stated in the instruction.
- Ensure each requirement is atomic, focusing on a single, indivisible aspect of the instruction.
- Format each requirement as a YES/NO question that can be used to verify if the requirement has been met.
Output Format
Provide the "evaluation requirements" only using numbered lists like the following format:
1. <evaluation requirement 1>
2. ...
Examples
### Given Instruction
Please explain Fermat’s Last Theorem within the context of an interview with a renowned mathematician, utilizing a FAQ template to address various aspects. Ensure the explanation is concise, limited to no more than three sentences.
### Evaluation Requirements
1. Does the response explain Fermat’s Last Theorem?
2. Is the explanation presented within the context of an interview with a renowned mathematician?
3. Does the response utilize a FAQ template to address various aspects?
4. Is the explanation concise?
5. Is the explanation limited to no more than three sentences?
Task Inputs & Outputs
Here’s the "given instruction". Let’s decompose evaluation requirements!
## Given Instruction
{given_instruction}
## Evaluation Requirements
Table 10: Prompt for Requirement Decomposition Step
The prompt used for requirement decomposition, as described in Section 3.1.3, is presented in Table 10. We use GPT-4o to decompose each core instruction in the Easy Set and each combined instruction in the Hard Set into multiple evaluation requirements for assessing LLM-generated responses. In our prompt, we explicitly instruct GPT-4o to generate a list of requirements that: (1) are strictly derived from the given instruction without inference, (2) are atomic and indivisible to enable fine-grained evaluation, and (3) are formatted as YES/NO questions for binary assessment. To ensure stable results, we set the maximum generation length to 512 tokens and use a temperature of 0.
One might argue that the brainstormed constraints could also serve as evaluation requirements. However, this does not eliminate the necessity of requirement decomposition. The need for decomposition arises not merely from the need to format them as questions, but due to the following reasons: (1) Evaluation requirements are essential for assessing the Easy Set instructions as well as the corresponding components within the Hard Set instructions. (2) The constraints generated by GPT-4o are sometimes compound statements encompassing multiple requirements. For example, a constraint like "Maintain a neutral but informative tone" actually consists of two distinct aspects: neutrality and informativeness. Furthermore, during the constraint combination step, these constraints may be rephrased or semantically altered, making explicit requirement decomposition necessary for accurate evaluation.
A.3.2 Requirement Categorization
Requirement Categorization Task
Your task is to categorize each item in the "evaluation requirements" into the most appropriate "requirement category" and "requirement direction".
The "evaluation requirements" are derived from the "given instruction," which provides the contextual background for each requirement.
Requirement Definitions and Categories
{{CONSTRAINT_TAXONOMY}}
Categorization Guidelines
Requirements should be categorized through a three-step evaluation process with category prioritization:
1. First, trace back each requirement to its original intent in the "given instruction", rather than focusing on how it’s phrased in the evaluation.
2. Second, evaluate if the requirement fits Numerical or Situation categories or not.
3. Third, if the requirement doesn’t fit Numerical or Situation categories, then consider Format, Style, or the lastly Content categories.
Output Format
Provide the "requirement categories" in a numbered list. For example:
1. <requirement category of 1st requirement> - <requirement direction of 1st requirement>
2. <requirement category of 2nd requirement> - <requirement direction of 2nd requirement>
Examples
### Given Instruction
Please explain Fermat’s Last Theorem within the context of an interview with a renowned mathematician, utilizing a FAQ template to address various aspects. Ensure the explanation is concise, limited to no more than three sentences.
### Evaluation Requirements
1. Does the response explain Fermat’s Last Theorem?
2. Is the explanation presented within the context of an interview with a renowned mathematician?
3. Does the response utilize a FAQ template to address various aspects?
4. Is the explanation concise?
5. Is the explanation limited to no more than three sentences?
### Requirements Categories
1. Content Requirements - Include Content
2. Situation Requirements - Create Situation/Environment
3. Format Requirements - Apply Template
4. Style Requirements - Specify Writing Style
5. Numerical Requirements - Limit Length Parameters
Task Inputs & Outputs
Here’s the "given instruction" and "evaluation requirements". Let’s categorize them properly!
## Given Instruction
{given_instruction}
## Evaluation Requirements
{evaluation_requirements}
## Requirements Categories
Table 11: Prompt for Requirement Categorization Step
The prompt used for requirement categorization in Section 3.1.3 is presented in Table 11. We provide GPT-4o with a list of evaluation requirements along with their instructions and ask it to map each requirement to the predefined constraint taxonomy for further analysis. The maximum generation length is set to 256 tokens, with a temperature of 0.
A.3.3 Requirement Assessment
Requirement Assessment Task
Task Introduction
Your task is to assess the evaluation requirements of each instruction against specific rules given in the criteria section. We will use evaluation requirements as criteria to assess instruction-following capabilities of Large Language Models (LLMs). These requirements provide fine-grained details on what should be included in model responses. While we will translate the instructions into different languages such as English, Chinese, Russian, Arabic, Hindi, Swahili, and more, the evaluation requirements will remain in English to ensure comparability across languages.
Input & Output Explanation
{{INPUT_OUTPUT_EXPLANANTION}}
Validation Criteria
Please check each instruction against the following criteria step-by-step:
Rule 1: Are the evaluation requirements explicitly stated in the instruction? [YES/NO]
Definition: Evaluation requirements should be explicitly stated in the instruction, without any need for inference or assumption. The requirements should be directly derived from the content of the instruction to ensure that the evaluation criteria are based on the user’s explicit intentions.
Positive Example:
   instruction: "What color goes best with teal?"
   evaluation_requirements: "1. content requirement - include content: Does the response identify a color that pairs well with teal?"
   Accepted Reason: The requirement is explicitly present in the instruction, asking for a color that pairs well with teal.
Negative Example:
   instruction: "What color goes best with teal?"
   evaluation_requirements: "2. content requirement - analyze components: Does the response explain why the suggested color goes well with teal?"
   Rejected Reason: The requirement is not explicitly present in the instruction, as it asks for an explanation rather than a color pairing.
Rule 2: Are the evaluation requirements complete? [YES/NO]
Definition: Evaluation requirements should encompass all explicitly stated elements from the instruction. Each component directly mentioned in the instruction should have a corresponding evaluation requirement. This ensures that the evaluation criteria cover all explicit instructions.
Positive Example:
   instruction: "As a journalist covering border news, could you provide information on whether the US border is open to Canada?"
   evaluation_requirements:
     - "1. situation requirement - assume role: Does the response adopt the perspective of a journalist covering border news?"
     - "2. content requirement - include content: Does the response provide information on whether the US border is open to Canada?"
   Accepted Reason: The requirements are complete, covering all explicitly stated elements from the instruction.
Negative Example:
   instruction: "As a journalist covering border news, could you provide information on whether the US border is open to Canada?"
   evaluation_requirements:
     - "1. content requirement - include content: Does the response provide information on whether the US border is open to Canada?"
   Rejected Reason: The requirements are incomplete, missing the situation requirement to adopt the journalist’s perspective.
Rule 3: Are the evaluation requirements atomic? [YES/NO]
Definition: Evaluation requirements should be atomic, focusing on a single, distinct aspect of the instruction. This means each requirement should be decomposed to the finest granularity where it can still be evaluated independently, without creating dependencies between requirements during evaluation.
Positive Example:
   instruction: "Create a short, concise summary of the paper based on its abstract."
   evaluation_requirements:
     - "1. content requirement - include content: Does the response provide a summary of the paper?"
     - "2. situation requirement - establish purpose: Is the summary based on the paper’s abstract?"
     - "3. style requirement - specify writing style: Is the summary short?"
     - "4. style requirement - specify writing style: Is the summary concise?"
   Accepted Reason: The requirements are atomic because each one focuses on a distinct aspect of the instruction.
Rule 4: Are the evaluation requirements properly categorized and directed? [YES/NO]
Definition: Evaluation requirements should be properly categorized and directed based on their primary function within the instruction, tracing back to the original intent rather than the phrasing. These categories will be used to analyze the instruction-following capabilities of LLMs from different perspectives. See {{CONSTRAINT_TAXONOMY}} for detailed definitions and categories.
Categorization Priority: Follow this order when categorizing requirements:
   1. Prioritize Numerical and Situation categories first, as they typically have the most explicit indicators.
   2. If the requirement does not fit into Numerical or Situation categories, then evaluate Format, Style, and finally Content categories.
Relation between Categorization and Direction:
   1. First, evaluate the requirement’s category based on its primary function.
   2. Next, evaluate the direction based on the instruction’s intent, ensuring the requirement aligns with the instruction’s purpose.
   3. Typically, if requirements are miscategorized, they will also be improperly directed.
Task Workflow
1. Read the instruction and evaluation_requirements fields.
2. Evaluate each evaluation_requirements against the validation criteria.
3. If feeling uncertain, you can (recommended from most to least encouraged):
   - Use web search or external LLMs to understand the instruction and requirements better.
   - Test with an LLM by:
     a. Input the instruction to get a response.
     b. Use the evaluation requirements to assess the response.
     c. Compare if your assessment aligns with the requirements’ intended evaluation.
4. Fill in the human_annotation_tags fields based on your assessment.
5. Add any additional comments or notes in the notes field if necessary.
Table 12: Human Annotation Guideline for Requirement Assessment Task
Here, we present the details of the requirement assessment introduced in Section 3.1.3. For each evaluation requirement, we assess its quality from four dimensions: explicitness, completeness, atomicity, and categorization. A detailed explanation of each dimension is provided in the annotation guideline (Table 12), resulting in five binary evaluation criteria for each requirement. The categorization dimension is assessed from two perspectives: (1) whether the requirement is correctly categorized and (2) whether it is assigned to the correct dimension. The completeness dimension is evaluated at the instruction level, while the other three dimensions are assessed at the requirement level.
To ensure the reliability of the assessment, two annotators independently annotated a randomly sampled subset, comprising 10% of the instructions from both sets (10 from the Easy Set and 50 from the Hard Set). We calculate both the proportion of high-quality requirements and the inter-annotator agreement on this subset. Disagreements were resolved using a conservative resolution strategy, similar to the approach described in Section A.2.4.
Criteria Ratio (%) Agreement (%)
Explicitness 99.8 99.9
Completeness 93.3 95.0
Atomicity 95.9 96.6
Categorization 94.8 96.3
Dimension 94.0 95.5
Table 13: Evaluation results for various criteria, presenting the Ratio of requirements or instructions (for completeness) that meet each criterion and the Agreement (%) between evaluators.
As shown in Table 13, almost all requirements adhere to the explicitness criterion, indicating that very few nonexistent requirements were generated. Additionally, the majority of evaluation requirements are rated as complete, atomic, and correctly categorized in both constraint category and dimension. These results demonstrate the high quality of the generated requirements. Furthermore, the inter-annotator agreement scores confirm the consistency and reliability of the annotation process.
A.4 Multilingual Expansion Details
Translation Validation Task
Your task is to evaluate and score the "translated instruction" in {translated_language} to determine if it maintains the essential requirements of the "original instruction," as detailed in the "evaluation requirements." Use both the "translated instruction" and the "back-translated instruction" to assess accuracy.
Your goal is to ensure all essential requirements from the original instruction are preserved in the translation, allowing for minor linguistic variations.
Input Definitions
Original Instruction: The original instruction that needs to be translated.
paper Translated Instruction:
   - The machine-translated version of the "original instruction" in {translated_language}.
   - It is the instruction needing to be evaluated for quality.
Back-translated Instruction:
   - The English machine-translated version of the "translated instruction".
   - It serves as an additional tool to verify the translation’s accuracy.
Evaluation Requirements:
   - Definition: YES/NO questions designed to assess whether a response follows the instruction. Each question corresponds to an explicit and atomic requirement from the original instruction.
   - Purpose: During evaluation, each question is used to locate the corresponding portions in the instruction that need to be evaluated in the translation.
   - Metadata: Information such as "content requirement - include content" indicates the category and direction of the requirement, aiding in identifying the portion to be evaluated.
   - Implicit Requirements: Occasionally, there may be implicit requirements not explicitly stated in the instruction. These should be identified and treated as exceptions (detailed below) to ensure they do not affect the overall validation.
Special Case: Implicit Requirements
Implicit requirements are requirements that are not directly present or reflected in the "original instruction" but may have been inferred or included in the "evaluation requirements" due to automated processing.
Do not evaluate the "translated instruction" for implicit requirements, as they do not exist in the original instruction and cannot be preserved in the translation. In this case, implicit requirements should be treated as exceptions and handled as follows:
1. Replace all portions (Original, Translated, and Back-translated) with the placeholder "[IMPLICIT_REQUIREMENT]".
2. Provide a concise observation explaining why the requirement is implicit.
3. Assign a score of "0/5" to indicate that the requirement is not applicable to the translation evaluation.
Scoring Scale
The scoring scale ranges from 1 to 5, with 3/5 as the acceptance threshold for translation quality.
- 5/5 (Perfect Preservation): Core requirements are preserved with complete semantic accuracy.
- 4/5 (Strong Preservation): Core requirements are maintained with very minor variations.
- 3/5 (Acceptable Preservation): Core requirements are maintained with some variations.
- 2/5 (Weak Preservation): Core requirements are partially distorted or unclear.
- 1/5 (Failed Preservation): Core requirements are significantly distorted or completely lost.
Remember to use a score of 0/5 as a special placeholder for implicit requirements, indicating their inapplicability.
Evaluation Guidelines
For each "evaluation requirement," follow these steps to assess the quality of the translation:
1. Portion Identification:
   - Extract the minimal relevant portion from "original instruction", "translated instruction", and "back-translated instruction" using the "evaluation requirements" as a guide.
   - If the instruction combines multiple requirements, isolate only the portion relevant to the current requirement to ensure focused evaluation.
   - For implicit requirements, replace all those 3 portions with the placeholder "[IMPLICIT_REQUIREMENT]".
2. Observation: Provide a concise observation focusing on:
   - How well the "translated portion" in "translated instruction" preserves the semantics of the "original portion" in "original instruction"?
   - Use the "back-translated portion" in "back-translated instruction" to verify the accuracy of the translation.
   - For implicit requirements, explain why the requirement is not applicable.
3. Scoring: Assign a score from 1 to 5 in the format "X/5" based on the scoring scale.
   - For implicit requirements, always assign a score of "0/5".
Output Format
Ensure your evaluation for each requirement is clearly structured as follows, without any additional information:
### Requirement {Number}: {Requirement Description}
#### Portions
- Original Portion: [Minimal relevant portion from "original instruction"]
- Translated Portion: [Minimal relevant portion from "translated instruction"]
- Back-translated Portion: [Minimal relevant portion from "back-translated instruction"]
#### Observation
{Provide a concise observation on semantic preservation and accuracy. For implicit requirements, explain why the requirement is implicit.}
#### Scoring
{Assign a score from 1 to 5 based on the scoring scale, formatted as "X/5". For implicit requirements, use "0/5".}
Task Inputs & Outputs
### Original Instruction
{original_instruction}
### Translated Instruction (in {translated_language})
{translated_instruction}
### Back-translated Instruction
{back_translated_instruction}
### Evaluation Requirements
{evaluation_requirements}
### Your Validations
Table 14: Prompt for Translation Validation Step
In this section, we detail the translation validation process introduced in Section 3.1.4. Since our annotators are not native speakers of Russian, Arabic, Hindi, or Swahili, we leverage Google Translate and GPT-4o to automatically assess translation quality. Our validation procedure follows a requirement-based evaluation framework, ensuring that each instruction and its translation are analyzed for both consistency and accuracy.
Specifically, for each instruction-translation pair, we prompt GPT-4o to identify corresponding segments in the English instruction and its translated counterpart, using predefined evaluation requirements as reference points. Additionally, we incorporate back-translation via Google Translate to mitigate potential biases, particularly for low-resource languages. The LLM annotator then provides a concise assessment of translation correctness and assigns a quality score on a 1-5 scale, where a score of 3 serves as the threshold for acceptable translation quality. The detailed prompt used for this evaluation is provided in Table 14.
Score Zh Ru Ar Hi Sw
5 93.46 95.09 93.26 94.31 89.53
4 5.30 3.29 4.38 4.30 7.60
3 0.93 1.01 0.97 0.81 1.63
2 0.27 0.54 0.85 0.39 0.97
1 0.04 0.08 0.54 0.19 0.27
Table 15: Distribution of translation quality scores across different languages (%).
To quantify translation quality, we analyze the distribution of evaluation requirement scores across different languages. The results, summarized in Table 15, show that the majority of translations receive high-quality ratings. Notably, the proportion of requirements scoring below 3 remains low across all languages, with Arabic exhibiting the highest at just 1.39%. These findings suggest that the translations are generally accurate and consistent with the original English instructions, and their inclusion in the benchmark does not significantly impact the overall evaluation results.
Appendix B Dataset Statistics
As illustrated in Figure 6, the distribution of evaluation requirements in the Easy Set is largely dominated by Content constraints, reflecting a lack of diversity in constraint types. However, despite being constructed based on the Easy Set, the Hard Set exhibits a more balanced distribution across all categories, indicating a diverse and well-balanced constraint augmentation process.
Figure 6: Requirement category distribution across the Easy Set and Hard Set.
Further analysis of the requirement count distribution across the Easy and Hard Sets is shown in Figure 7. While both sets follow a similar uniform distribution, the Hard Set exhibits a broader range of requirement counts, with a pronounced peak at 5 requirements, compared to 2 in the Easy Set. This distribution highlights the greater diversity and complexity of the Hard Set, making it a more challenging evaluation benchmark for LLMs.
Figure 7: Requirement count distribution across the Easy Set and Hard Set.
Appendix C Evaluation Protocol Details
Instruction Following Evaluation Task
Your task is to evaluate whether the "model response" follows the "English instruction" by checking if it satisfies all "evaluation requirements".
Your goal is to assess the instruction following quality of the "model response" using binary (YES/NO) decisions for each evaluation requirement.
Input Definitions
• English Instruction:
   - The English instruction given to a Large Language Model (LLM).
   - It also serves as the source for deriving evaluation requirements.
• Model Response:
   - The LLM’s response to the English instruction.
   - It will be evaluated against each requirement.
• Evaluation Requirements:
   - The YES/NO questions designed to assess whether the response follows the instruction. Each question corresponds to an explicit and atomic requirement from the English instruction.
   - Additional metadata (e.g., "content requirement - include content") provides context about the type and focus of the requirement. Use this metadata to identify the specific aspect of the response that needs to be evaluated.
Evaluation Rules
For each evaluation requirement, provide a strict YES/NO decision based on the following principles:
• YES: Select "YES" only if the response fully satisfies the requirement without any errors, omissions, ambiguities, or deviations, even minor ones.
• NO: Select "NO" if the response fails to satisfy the requirement, contains inaccuracies, omits relevant details, or introduces ambiguity.
This strict evaluation means that even if the response is mostly correct, partially correct, or correct under certain conditions, it should still be evaluated as "NO".
Output Format
Provide a structured evaluation for each requirement in order based on the following format, without any additional information.
## Requirement {Number}: {Requirement Description}
### Observation
 [Provide a concise observation on how the "model response" satisfies the requirement.]
### Decision
 [Output your evaluation decision (YES/NO) for the requirement only.]
Task Inputs & Outputs
## English Instruction
{english_instruction}
## Model Response
{model_response}
## Evaluation Requirements
{evaluation_requirements}
## Your Evaluation
Table 16: Prompt for the Instruction Following Evaluation Task on English Instructions.
Instruction Following Evaluation Task
Your task is to evaluate whether the "model response" follows the "{translated_language} instruction" by checking if it satisfies all "evaluation requirements".
Your goal is to assess the instruction following quality of the "model response" using binary (YES/NO) decisions for each evaluation requirement.
Input Definitions
• {translated_language} Instruction:
   - The {translated_language} instruction given to a Large Language Model (LLM).
   - The directive part is translated from the English instruction, while other parts remain unchanged intentionally.
• Model Response:
   - The LLM’s response to the {translated_language} instruction.
   - It will be evaluated against each requirement.
• English Instruction:
   - The original English instruction from which the {translated_language} instruction is translated. Unchanged parts are omitted for brevity.
   - It serves as the source for deriving evaluation requirements and helps in understanding the {translated_language} instruction.
• Evaluation Requirements:
   - The YES/NO questions designed to assess whether the response follows the instruction. Each question corresponds to an explicit and atomic requirement from the English instruction.
   - Additional metadata (e.g., "content requirement - include content") provides context about the type and focus of the requirement. Use this metadata to identify the specific aspect of the response that needs to be evaluated.
Evaluation Rules
For each evaluation requirement, provide a strict YES/NO decision based on the following principles:
• YES: Select "YES" only if the response fully satisfies the requirement without any errors, omissions, ambiguities, or deviations, even minor ones.
• NO: Select "NO" if the response fails to satisfy the requirement, contains inaccuracies, omits relevant details, or introduces ambiguity.
This strict evaluation means that even if the response is mostly correct, partially correct, or correct under certain conditions, it should still be evaluated as "NO".
Output Format
Provide a structured evaluation for each requirement in order based on the following format, without any additional information.
## Requirement {Number}: {Requirement Description}
### Observation
 [Provide a concise observation on how the "model response" satisfies the requirement.]
### Decision
 [Output your evaluation decision (YES/NO) for the requirement only.]
Task Inputs & Outputs
## {translated_language} Instruction
{translated_instruction}
## Model Response
{model_response}
## English Instruction
{english_instruction}
## Evaluation Requirements
{evaluation_requirements}
## Your Evaluation
Table 17: Prompt for the Instruction Following Evaluation Task on Translated Instructions.
The prompt used in the evaluation framework (Section 3.3) is provided in Table 16 for English instructions and Table 17 for other languages. For English instructions, GPT-4o receives the instruction, the corresponding response, and the evaluation requirements to assess the response’s adherence to the given instruction. The model first provides observations on the response, followed by a binary YES/NO decision for each requirement, indicating whether the response meets it. For instructions in other languages, we follow a similar procedure, providing GPT-4o with the translated instruction, the corresponding response, the original English instruction, and the evaluation requirements. To ensure stable results, we set the maximum generation length to 2048 tokens and use a temperature of 0.
Appendix D Experiment Details
D.1 Automatic Evaluation
The official API names of the models used in Section 4.1 are: gpt-4o-2024-11-20, gemini-2.0-flash-001 and claude-3-5-sonnet-20241022. The open-source models and their repositories on Hugging Face [^6] https://huggingface.co/models are: THUDM/glm-4-9b-chat, meta-llama/Llama-3.1-8B-Instruct, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-7B-Instruct, Qwen/Qwen2.5-14B-Instruct, and Qwen/Qwen2.5-72B-Instruct. Evaluating an LLM on the entire XIFBench dataset using gpt-4o-2024-08-06 (without batching) costs approximately \$24.
D.2 Agreement Evaluation
Instruction Following Evaluation Assessment
Task Introduction
Your task is to assess the decisions made by a GPT-4o based evaluator and determine whether you agree with them. This multilingual instruction-following benchmark covers six languages: English (en), Chinese (zh-cn), Russian (ru), Arabic (ar), Hindi (hi), and Swahili (sw). Your evaluations will help us measure the agreement of GPT-4o based evaluators with humans across different languages.
Input & Output Explanation
You will receive a TXT file and a JSON file, each containing the prompts, responses and GPT-4o’s evaluations. Each item in both files shares the same unique ID.
{{INPUT_OUTPUT_EXPLANATION}}
Output Explanation
You need to review each decision made by the GPT-4o evaluator and indicate whether you agree with their assessment. Record your agreement in the instruction_following_evaluation section of the JSON file under the following fields: - do_agree: Set to true if you agree with the evaluator’s decision. Set false if you disagree.
   - note: (Optional) Provide a brief reason if you disagree with the evaluator’s decision. Important Notes:
   - The initial true or false values in do_agree are placeholders – they should not influence your judgment.
   - While the note field is optional, it is strongly recommended when you disagree, as it helps us to analyze disagreement cases.
Task Guidelines
The satisfaction of each requirement in decision should be evaluated with a strict binary approach (true or false). Below are the principles of judgment for both the GPT-4o evaluator and you: - true: Select true only if the response fully satisfies the requirement without any errors, omissions, ambiguities, or deviations, no matter how minor.
   - false: Select false if the response fails to satisfy the requirement, contains inaccuracies, omits relevant details, or introduces ambiguity. This strict evaluation means that even if the response is mostly correct, partially correct, or correct under certain conditions, it should still be evaluated as false.
Task Examples
The following examples illustrate how strict GPT-4o and you should be for assessing instruction-following quality. These are not exhaustive but should guide your judgment. - content requirement - connect elements: Does the response connect the meaning of life to human happiness?
      Select false if the response relates the meaning of life to human happiness improperly or connects elements ambiguously.
   - style requirement - specify writing style: Is the response written in an academic style?
      Select false if the response contains informal language or lacks academic rigor.
   - situation requirement - assume role: Does the response adopt the perspective of a historian?
      Select false if the response does not clearly adopt a historian’s perspective or rarely implies a historian’s viewpoint.
   - situation requirement - establish purpose: Does the response aim to persuade the reader?
      Select false if the response fails to persuade the reader or persuades the reader ambiguously.
   - format requirement - specify output format/pattern: Is the response formatted as a checklist?
      Select false if the response lacks clear checkboxes, bullet points or numbering.
   - numerical requirement - limit length parameters: Does the response have exactly five sentences?
      Select false if the response has more or fewer than five sentences.
Task Workflow
1. Read the original_instruction in the JSON file to understand the task requirements.
2. For each requirement in instruction_following_evaluation, follow these steps:
   - Read the requirement and observation to understand the context.
   - Read the [PROMPT] and [RESPONSE] sections in the TXT file to examine the actual prompt and response.
   - Use Google Translate or other tools if needed to understand the [PROMPT] and [RESPONSE] content.
   - Determine if the [RESPONSE] fully meets the requirement.
   - The observation can help navigate the response, but do not rely on it as evidence for agreement or disagreement, as it may not always accurately reflect the response.
3. Update the do_agree field based on your assessment for each requirement.
4. If you disagree with the evaluator’s decision, provide a brief reason in the note field (if possible).
Table 18: Human Annotation Guideline for Instruction Following Evaluation Assessment Task
In this section, we describe the process of evaluating the agreement between GPT-4o and human annotations. We randomly sample 3 English instructions from the Easy Set and 17 from the Hard Set for representation. Their translations are also included to ensure cross-linguistic comparability, resulting in a total of 120 instructions. To assess diverse response quality, we collect outputs from three representative models: Gemini-2.0-Flash, Qwen-2.5-72B, and GLM-4-9B. This yields 360 instruction-response pairs for evaluation.
Two annotators then evaluate whether GPT-4o’s assessments align with human judgments for each evaluation criterion, following the guidelines in Table 18. Additionally, GPT-4o’s generated observations during decision-making are provided to help annotators identify relevant response segments. The final ground truth is established by reversing GPT-4o’s decisions in case of disagreements. Finally, we calculate the agreement rates between GPT-4o and human annotations to assess the reliability of our evaluation protocol and direct scoring.